OpenStudy (anonymous):

Randy and Trey take turns cleaning offices on the weekends. It takes Randy at most 4 hours to clean the offices. It takes Trey at most 6 hours to clean the offices. What is the greatest amount of time it would take them to clean the offices together?

OpenStudy (kropot72):

In one hour Randy cleans at least 1/4 of the offices and in one hour Trey cleans at least 1/6 of the offices. Working together they would clean at least 1/4 + 1/6 = 5/12 of the offices. Let x be the greatest amount of time it would take them to clean the offices together. Then we can write the following equation: \[\frac{5x}{12}=1\] Now you just need to solve the equation to find x. Can you do that?

OpenStudy (anonymous):

2.4 I got the answer thanks

OpenStudy (kropot72):

You're welcome :)